# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v2

@@ 3.1 Finite-time geodesic structure
Inverting eq.˜15, its solution reads:
@@
Equivalently, using eq.˜16, the finite-time optimal protocol for $t\in(0,\tau)$ can be expressed using the known control parameters $\mathcal{G}(A_{\alpha}^{i})$ and $\mathcal{G}(A_{\alpha}^{f})$ as
@@
Equation˜18 is the optimal finite-time transport map. It is equivalent to substituting $t/\tau\to t/(\tau+2)$, $1\to{(1+\tau)}/{(2+\tau)}$, and $0\to{1}/{(2+\tau)}$ in eq.˜7. Physically, optimal finite-time driving is equivalent to total driving time $\tau+2$ with initial and final times $t=1$ and $t=\tau+1$, respectively, leading to jumps in the optimal protocol at the endpoints. This reveals the structural geometric similarities between the finite-time optimal driving and the slow-driving process.
Using eq.˜18, the amplitude of kinks is quantified as $\Delta_{0+}\mathcal{G}_{\tau}(A_{\alpha})=\Delta_{\tau-}\mathcal{G}_{\tau}(A_{\alpha})=\frac{1}{2+\tau}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)$. Hence, kinks are of equal amplitude in geodesic space, which physically corresponds to equal distribution of the thermodynamic cost associated with kinks. Upon converting back to the control parameter space $A_{\alpha}$ using $\mathcal{G}^{-1}$, kinks have different amplitudes. The higher-mass endpoint has a smaller jump size compared to the lower-mass endpoint. The amplitude of kinks scales as $1/(2+\tau)$, vanishing in the quasistatic limit $\tau\to\infty$. For the fast-driving limit $\tau\to 0$, the optimal protocol follows midpoint interpolation in geodesic space,
@@
Inverting back to the control parameter space gives $A_{\alpha}=\mathcal{G}^{-1}\left(\frac{1}{2}\mathcal{G}(A_{\alpha}^{i})+\frac{1}{2}\mathcal{G}(A_{\alpha}^{f})\right)$ in the limit $\tau\to 0$. $\mathcal{G}_{\tau}(A_{\alpha})$ is plotted in fig.˜1(e) for different $\tau$. Due to $\mathcal{G}(A_{\alpha})>A_{\alpha}$, endpoint jumps are amplified for fEQ systems.
Using the finite-time optimal transport map eq.˜18, the optimal bulk and boundary EP are
@@
Equation˜20 reveals the fundamental trade-off between bulk and boundary EP in optimal finite-time processes. In particular, $\Sigma_{bulk}^{*}\propto\tau$ and $\Sigma_{bnd}^{*}\propto 1$, implying that finite-time optimal protocols with kinks are a physical manifestation of the optimization interplay between them. The large-$\tau$ and small-$\tau$ regimes correspond to the domination of $\Sigma_{bulk}^{*}$ and $\Sigma_{bnd}^{*}$, respectively, for the thermodynamic optimization. The total optimal driving EP $\Sigma_{\tau}^{*}=\Sigma_{bulk}^{*}+\Sigma_{bnd}^{*}$ is,
@@
Compared to eq.˜8, $\Sigma_{\tau}^{*}$ is finite in the $\tau\to 0$ limit and smaller than $\Sigma_{qs}^{*}$. They satisfy the non-equilibrium scaling relations $\Sigma_{\tau}^{*}\propto 1/(2+\tau)$ and $\Sigma_{qs}^{*}\propto 1/\tau$, respectively. Importantly, eq.˜21 imposes an upper bound on the driving dissipation $\Sigma_{\tau}^{*}\leq\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2}/4$, saturated in the $\tau\to 0$ limit. An analogy of an upper bound for the driving dissipation is missing for the slow-driving processes.
@@
Thermodynamically, finite-time optimal driving is equivalent to slow driving with driving time $\tau+2$ instead of $\tau$. This interpretation is consistent with the dynamic mapping between slow-driving and finite-time optimal protocols elucidated using eqs.˜7 and 18. The short-time underestimation from $\Sigma_{\tau}\propto 1/\tau$ scaling has been experimentally observed [30]. In conclusion, we analytically solve the GFTOC problem by revealing its exact connection to its slow-driving counterpart.
Figure 2: (a) The finite-time optimal protocol $\mathcal{G}_{\tau}(F_{\alpha})\to t$ is plotted for the different values of $\tau=\{4,1,0.1\}$ with the same initial and final value condition (shown by the dotted blue lines) with $F_{\alpha}^{i}=2$ and $F_{\alpha}^{f}=5$. Its slow-driving counterpart is denoted by ‘qs’. The vanishing excess affinity is considered because the optimal control of the housekeeping EPR is being solved. Where, eqs.˜7, 18 and 9 are used for the plot and the scaled (with $\tau$) driving time ($t/\tau$) is used. Kinks are of equal amplitude in the geodesic space. (b) The corresponding finite-time non-conservative affinity is obtained using the fEQ geodesic $\mathcal{G}^{fEQ}(F_{\alpha})$ in eq.˜9, which gives the mapping $F_{\alpha}\to t$ in scaled (with $\tau$) driving time ($t/\tau$). Kinks are of unequal amplitude in the affinity space, such that a higher mass point exhibits a smaller amplitude of kinks. (c) The corresponding mapping in $\mathcal{G}_{\tau}(F_{\alpha})\to F_{\alpha}$ space, the initial and final points are denoted by $\star$ and $\blacklozenge$, respectively. Due to the finite driving time constraint, the finite-time optimal protocol eq.˜18 traverses a part of the slow-driving geodesic eq.˜7, such that the finite-time speed limit eq.˜22 is restored, analogous to the slow-driving speed limit eq.˜13. The choice of discontinuous jumps in the finite-time driving protocol is also constrained by equally distributing the thermodynamic cost ($\Sigma_{bnd}^{*}$ in eq.˜20) between the initial and final endpoints, which generates a ‘Thermodynamic shock’ at the initial and final time.
3.2 The physical interpretation of kinks and the role of boundary conditions
@@
We elaborate on the physical interpretation of kinks within our GFTOC framework. In the small driving-time limit, the fundamental principle of ST—the timescale separation between the environment and system degrees of freedom—is violated. This principle normally ensures that the system instantaneously relaxes to the control parameters imposed by the environment. Simultaneously, the boundary conditions enforce a constraint of reaching the final control parameters within the given finite time $\tau$ from the specified initial control parameters. Hence, for small $\tau$, due to violation of the speed-limit bound eq.˜13, the environment cannot instantaneously impose the required control parameters on the system.
@@
Kinks solve this issue, as they reduce the geodesic distance for the finite-time driving process, mathematically expressed as $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$. This trade-off between the rush to reach the final point along the geodesic and the inability to do so within time $\tau$ produces kinks. The thermodynamic cost of each kink is $\Sigma_{bnd}^{*}/2$. The system undergoes a ‘thermodynamic shock’ at the initial and final points, generating a spontaneous discontinuous change in the control parameters, which induces a global transition of the system’s state. Within our framework, $\Sigma_{bnd}^{*}$ and $\Sigma_{bulk}^{*}$ can be interpreted analogously to heat and work, respectively: heat corresponds to the instantaneous discrete cost of a jump, and work to the continuous driving of control parameters. To our knowledge, this provides a novel thermodynamic understanding of the heat–work dual manifestation of the driving EP in finite-time optimal driven far-from equilibrium processes.
3.3 Finite-time speed limit
Using $\Sigma_{bulk}^{*}$ from eq.˜20 and eqs.˜13 and 16, one finds
@@ 3.3 Finite-time speed limit
This shows that the selection of endpoint jump values $\mathcal{G}(A_{\alpha}^{i*})$ and $\mathcal{G}(A_{\alpha}^{f*})$ is constrained to restore the speed-limit bound eq.˜13 for the finite-time optimal driving process. Physically, this implies that the allowed endpoint jumps are minimal, just sufficient to restore the finite-time speed-limit constraint for any given $\tau$. This highlights a fundamental symmetry: kinks are a non-equilibrium mechanism that circumvents the dynamic constraint of limited driving time quantified by the speed-limit, representing the dynamic counterpart of a ‘thermodynamic shock’.
@@
The upper bound on the total amplitude of kinks, $|\Delta_{0^{+}}\mathcal{G}_{\tau}(A_{\alpha})|+|\Delta_{\tau^{-}}\mathcal{G}_{\tau}(A_{\alpha})|$, a boundary property, can be obtained using the thermodynamic cost of driving and the distance covered in control parameter space (a bulk property). The equal thermodynamic cost $\Sigma_{bnd}^{*}$ for each kink results from minimizing $\Sigma_{bnd}^{*}$ for a constrained total amplitude of kinks. Hence, kinks arise from finite driving-time constraints that may prevent reaching the final state along the geodesic. The finite-time optimal protocol restores eq.˜13 while minimizing the thermodynamic cost of both kinks. Importantly, the slow-driving assumption is broken for any finite $\tau$ and does not require $\tau<\tau_{qs}^{*}$. A continuous change of $\tau$ therefore induces a smooth transition from eqs.˜7 and 8 to eqs.˜17 and 21, with $\tau$ as the relevant control parameter. Figure˜2 pictorially summarizes the underlying geometric structure of GFTOC.
3.4 Time-reversal symmetry of geodesics
@@
Equations˜7 and 18 satisfy time-reversal symmetry, as they are invariant under the transformation $\mathcal{G}(A_{\alpha}^{i})\to\mathcal{G}(A_{\alpha}^{f})$, $\mathcal{G}(A_{\alpha}^{f})\to\mathcal{G}(A_{\alpha}^{i})$, and $t\to\tau-t$. This symmetry arises fundamentally from the geodesic structure of the slow-driving process. The symmetric amplitude of kinks extends this symmetry to the finite-time geodesic eq.˜17.
3.5 The experimental inference of the geodesic and
the non-equilibrium scaling relation
@@ 3.5 The experimental inference of the geodesic and
GFTOC framework implies that the minimum driving EP is obtained by computing the geodesic. If experimental setups can measure the optimal driving EP (thermodynamically constrained to support a minimum dissipation for the driving process), one can pose the inverse problem of reconstructing the geodesic for given initial and final control parameters. The non-equilibrium scaling relation eq.˜21 can be used to infer the analytical shape of $\mathcal{G}(A_{\alpha})$. For observed $(\Sigma_{\tau}^{*},\tau)$ data at fixed $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$, the collapse of data yields $\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})=\sqrt{2(2+\tau)\Sigma_{\tau}^{*}}$. Repeating this procedure for different $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$ reconstructs $\mathcal{G}(A_{\alpha})$ experimentally or numerically. The geodesic shape informs about the non-equilibrium character of the process and the contribution of EP, since it has different analytical dependence on the cEQ and fEQ regimes, as discussed before in section˜2.2 and later in section˜5.
@@ 3.6 Comparison and differences to model-specific known results
Finite-time optimal protocols with kinks were analytically computed for a particle in a harmonic trap [7] and recently unified using OMTT [8]. These model-specific results differ from the generalized finite-time framework in three aspects: (i) they optimize the free energy $\psi_{E}$ using control parameter $E$, (ii) the boundary conditions used to obtain kinks differ, and (iii) the thermodynamic cost of kinks is undefined in Ref.[7, 8], leading to the physical interpretation of kinks as a boundary artefact.
@@
We shortly summarize the procedure outlined in Ref.[8]. The boundary conditions used in [7, 8] are equivalent to fixing $A_{\alpha}^{i*}=A_{i}$ and computing $A_{\alpha}^{*int}\in(A_{\alpha}^{i},A_{\alpha}^{f})$ that minimises the bulk driving EP under the final value constraint. Thus, $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ is termed as the feasible driving protocol for the given finite-time $\tau$. Then, $A_{\alpha}^{*int}$ is used to compute $A_{\alpha}^{*i}$ and $A_{\alpha}^{*f}$ by introducing counter-adiabatic driving terms that shifts the initially computed driving protocol $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ to $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ in driving time $\tau$, and it generates kinks of amplitude $A_{\alpha}^{i}-A_{\alpha}^{*i}$ and $A_{\alpha}^{f}-A_{\alpha}^{f*}$ in the optimal protocol. This implies that the computation of the initial geodesic for the finite-time process is based on the assumption of the absence of kinks at the initial time $t=0$. However, the finite-time optimal protocol $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ is obtained by shifting this protocol (counter-adiabatic driving) to obtain the kink at both $t=0^{+}$ and $t=\tau^{-}$.
@@
This contradicts its formulation of constrained and unconstrained optimization for the initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, respectively. Hence, mathematically, this is interpreted as a boundary term artefact because of the inability of assigning a thermodynamic cost to kinks within their framework. In contrast, within MinAP formalism, a thermodynamic cost is inherently assigned for kinks, and hence unconstrained optimization is implemented for both initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, signified by a double optimization problem in eq.˜14. This implies our optimization framework uses minimum assumptions, equivalently assuming the least information about optimal protocols. This allows us to implement an unconstrained optimization problem in eq.˜14 that searches/accesses a wider control parameter space for the optimal protocol.
@@
These differences lead to rather drastic physical results. For example, the short-time optimal protocol eq.˜19 is midpoint interpolation in the geodesic space $\mathcal{G}(A_{\alpha})$ and not necessarily in control parameter space $A_{\alpha}$ unless $\mathcal{G}(A_{\alpha})=A_{\alpha}$. Hence, even for short driving time $\tau$, the finite-time optimal protocol eq.˜19 respects the Riemannian geometric structure in control parameter space. In contrast, the short-time protocols obtained in Ref. [7, 8] follow a midpoint interpolation in control parameter space. This difference is irrelevant in the case of the optimal control of the trap centre, since $\mathcal{G}(A_{\alpha})=A_{\alpha}$ for the control of the trap centre and the short-time optimal protocols obtained in Ref. [70, 71]. However, $\mathcal{G}(A_{\alpha})\neq A_{\alpha}$ for the optimal control of the trap stiffness [17], giving divergent analytical results compared to Ref. [7, 8, 70, 71]. Subsequently, we will revisit this issue in section˜5.3 with the example of a stochastic particle in a harmonic trap and reconcile the discrepancies between OMTT [39, 40, 51] and thermodynamic geometry [17].

### v3

@@ 3.1 Finite-time geodesic structure
Inverting eq.˜18, its solution reads:
@@
Equivalently, using eq.˜19, the finite-time optimal protocol for $t\in(0,\tau)$ can be expressed using the known control parameters $\mathcal{G}(A_{\alpha}^{i})$ and $\mathcal{G}(A_{\alpha}^{f})$ as
@@
Equation˜21 is the optimal finite-time transport map. It is equivalent to substituting $t/\tau\to t/(\tau+2)$, $1\to{(1+\tau)}/{(2+\tau)}$, and $0\to{1}/{(2+\tau)}$ in eq.˜8. Physically, optimal finite-time driving is equivalent to total driving time $\tau+2$ with initial and final times $t=1$ and $t=\tau+1$, respectively, leading to jumps in the optimal protocol at the endpoints. This reveals the structural geometric similarities between the finite-time optimal driving and the slow-driving process. Using eq.˜21, the amplitude of kinks is,
$\begin{split}\Delta_{0+}\mathcal{G}_{\tau}(A_{\alpha})=\Delta_{\tau-}\mathcal{G}_{\tau}(A_{\alpha})=\frac{1}{2+\tau}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right).\end{split}$ (22)
Hence, kinks are of equal amplitude in geodesic space, which physically corresponds to equal distribution of the thermodynamic cost associated with kinks. Upon converting back to the control parameter space $A_{\alpha}$ using $\mathcal{G}^{-1}$, kinks have different amplitudes. The higher-mass endpoint has a smaller jump size compared to the lower-mass endpoint. The amplitude of kinks scales as $1/(2+\tau)$, vanishing in the quasistatic limit $\tau\to\infty$. For the fast-driving limit $\tau\to 0$, the optimal protocol follows midpoint interpolation in geodesic space,
@@
Inverting back to the control parameter space gives $A_{\alpha}=\mathcal{G}^{-1}\left(\frac{1}{2}\mathcal{G}(A_{\alpha}^{i})+\frac{1}{2}\mathcal{G}(A_{\alpha}^{f})\right)$ in the limit $\tau\to 0$. $\mathcal{G}_{\tau}(A_{\alpha})$ is plotted in fig.˜2(a) for different $\tau$. Due to $\mathcal{G}(A_{\alpha})>A_{\alpha}$, endpoint jumps are amplified for fEQ systems.
Using the finite-time optimal transport map eq.˜21, the optimal bulk and boundary EP are
@@
Equation˜24 reveals the fundamental trade-off between bulk and boundary EP in optimal finite-time processes. In particular, $\Sigma_{bulk}^{*}\propto\tau$ and $\Sigma_{bnd}^{*}\propto 1$, implying that finite-time optimal protocols with kinks are a physical manifestation of the optimization interplay between them. The large-$\tau$ and small-$\tau$ regimes correspond to the domination of $\Sigma_{bulk}^{*}$ and $\Sigma_{bnd}^{*}$, respectively, for the thermodynamic optimization. The total optimal driving EP is, $\Sigma_{\tau}^{*}=\Sigma_{bulk}^{*}+\Sigma_{bnd}^{*}$,
@@
Compared to eq.˜9, $\Sigma_{\tau}^{*}$ is finite in the $\tau\to 0$ limit and smaller than $\Sigma_{qs}^{*}$. They satisfy the non-equilibrium scaling relations $\Sigma_{\tau}^{*}\propto 1/(2+\tau)$ and $\Sigma_{qs}^{*}\propto 1/\tau$, respectively. Importantly, eq.˜25 imposes an upper bound on the driving dissipation,
$\begin{split}\Sigma_{\tau}^{*}\leq\frac{1}{4}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2},\end{split}$ (26)
which is saturated in the $\tau\to 0$ limit. An analogy of an upper bound eq.˜26 is missing for the slow-driving processes.
@@
Thermodynamically, finite-time optimal driving is equivalent to slow driving with driving time $\tau+2$ instead of $\tau$. This interpretation is consistent with the dynamic mapping between slow-driving and finite-time optimal protocols elucidated using eqs.˜8 and 21. The short-time underestimation from $\Sigma_{\tau}\propto 1/\tau$ scaling has been experimentally observed [37]. In conclusion, we analytically solve the GFTOC problem by revealing its exact geometric connection to its slow-driving counterpart.
3.2 The physical interpretation of kinks and
the role of boundary conditions
@@ 3.2 The physical interpretation of kinks and
We elaborate on the physical interpretation of kinks within our GFTOC framework. In the small driving-time limit, the fundamental principle of ST—the timescale separation between the environment and system degrees of freedom—is violated. This principle normally ensures that the system instantaneously relaxes to the control parameters imposed by the environment. Simultaneously, the boundary conditions enforce a constraint of reaching the final control parameters within the given finite time $\tau$ from the specified initial control parameters. Hence, for small $\tau$, due to violation of the speed-limit bound eq.˜15, the environment cannot instantaneously impose the required control parameters on the system.
@@
Kinks solve this issue, as they reduce the geodesic distance for the finite-time driving process, mathematically expressed as $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$. This trade-off between the rush to reach the final point along the geodesic and the inability to do so within time $\tau$ produces kinks. The thermodynamic cost of each kink is $\Sigma_{bnd}^{*}/2$. The system undergoes a ‘thermodynamic shock’ at the initial and final points, generating a spontaneous discontinuous change in the control parameters, which induces a global transition of the system’s state. Within our framework, $\Sigma_{bnd}^{*}$ and $\Sigma_{bulk}^{*}$ can be interpreted analogously to heat and work, respectively: heat corresponds to the instantaneous discrete entropy production cost of a jump, and work to the continuous driving of control parameters. To our knowledge, this provides a novel thermodynamic understanding of the heat–work dual manifestation of the driving EP in finite-time optimal driven far-from equilibrium processes.
3.3 Finite-time geometric speed limit
Using $\Sigma_{bulk}^{*}$ from eq.˜24 and eqs.˜15 and 19, one finds
@@ 3.3 Finite-time geometric speed limit
This shows that the selection of endpoint jump values $\mathcal{G}(A_{\alpha}^{i*})$ and $\mathcal{G}(A_{\alpha}^{f*})$ is constrained to restore the speed-limit bound eq.˜15 for the finite-time optimal driving process. Physically, this implies that the allowed endpoint jumps are minimal, just sufficient to restore the finite-time speed-limit constraint for any given $\tau$. This highlights a fundamental symmetry: kinks are a non-equilibrium mechanism that circumvents the dynamic constraint of limited driving time quantified by the speed-limit, representing the dynamic counterpart of a ‘thermodynamic shock’.
@@
The upper bound on the total amplitude of kinks, $|\Delta_{0^{+}}\mathcal{G}_{\tau}(A_{\alpha})|+|\Delta_{\tau^{-}}\mathcal{G}_{\tau}(A_{\alpha})|$, a boundary property, can be obtained using the thermodynamic cost of driving and the distance covered in control parameter space (a bulk property). The equal thermodynamic cost $\Sigma_{bnd}^{*}$ for each kink results from minimizing $\Sigma_{bnd}^{*}$ for a constrained total amplitude of kinks. Hence, kinks arise from finite driving-time constraints that may prevent reaching the final state along the geodesic. The finite-time optimal protocol restores eq.˜15 while minimizing the thermodynamic cost of both kinks. Importantly, the slow-driving assumption is broken for any finite $\tau$ and does not require $\tau<\tau_{qs}^{*}$. A continuous change of $\tau$ therefore induces a smooth transition from eqs.˜8 and 9 to eqs.˜20 and 25, with $\tau$ as the relevant control parameter. In conclusion, this mechanism is to be interpreted as a ‘geometric shortcut’. Figure˜2 pictorially summarizes the underlying geometric structure of GFTOC.
3.4 Key features of geodesic
@@
Time-reversal symmetry. —  Equations˜8 and 21 satisfy time-reversal symmetry, as they are invariant under the transformation $\mathcal{G}(A_{\alpha}^{i})\to\mathcal{G}(A_{\alpha}^{f})$, $\mathcal{G}(A_{\alpha}^{f})\to\mathcal{G}(A_{\alpha}^{i})$, and $t\to\tau-t$. This symmetry arises fundamentally from the geodesic structure of the slow-driving process. The symmetric amplitude of kinks extends this symmetry to the finite-time geodesic eq.˜20.
Figure 2: (a) The finite-time optimal protocol $\mathcal{G}_{\tau}(F_{\alpha})\to t$ is plotted for the different values of $\tau=\{4,1,0.1\}$ with the same initial and final value condition (shown by the dotted blue lines) with $F_{\alpha}^{i}=2$ and $F_{\alpha}^{f}=5$. Its slow-driving counterpart is denoted by ‘qs’. As formulated later in section˜5.5, the vanishing excess affinity is considered because the optimal control of the housekeeping EPR is being solved. Where, eqs.˜8, 21 and 10 are used for the plot and the scaled (with $\tau$) driving time ($t/\tau$) is used. Kinks are of equal amplitude in the geodesic space. (b) The corresponding finite-time non-conservative affinity is obtained using the fEQ geodesic $\mathcal{G}^{fEQ}(F_{\alpha})$ in eq.˜10, which gives the mapping $F_{\alpha}\to t$ in scaled (with $\tau$) driving time ($t/\tau$). Kinks are of unequal amplitude in the affinity space, such that a higher mass point exhibits a smaller amplitude of kinks. (c) The corresponding mapping in $\mathcal{G}_{\tau}(F_{\alpha})\to F_{\alpha}$ space, the initial and final points are denoted by $\star$ and $\blacklozenge$, respectively. Due to the finite driving time constraint, the finite-time optimal protocol eq.˜21 traverses a part of the slow-driving geodesic eq.˜8, a ‘geometric shortcut’, such that the finite-time geometric speed limit eq.˜27 is restored, analogous to the slow-driving geometric speed limit eq.˜15. The choice of discontinuous jumps in the finite-time driving protocol is also constrained by equally distributing the thermodynamic cost ($\Sigma_{bnd}^{*}$ in eq.˜24) between the initial and final endpoints, which generates a ‘Thermodynamic shock’ at the initial and final time.
@@
Experimental inference and non-equilibrium scaling relation. —  GFTOC framework implies that the minimum driving EP is obtained by computing the geodesic. If experimental setups can measure the optimal driving EP (thermodynamically constrained to support a minimum dissipation for the driving process), one can pose the inverse problem of reconstructing the geodesic for given initial and final control parameters. The non-equilibrium scaling relation eq.˜25 can be used to infer the analytical shape of $\mathcal{G}(A_{\alpha})$. For the observed set of data $\{(\Sigma_{\tau}^{*},\tau)\}$ at fixed $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$, the collapse of the data yields $\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})=\sqrt{2(2+\tau)\Sigma_{\tau}^{*}}$. Repeating this procedure for different $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$ reconstructs $\mathcal{G}(A_{\alpha})$ experimentally or numerically. The geodesic shape informs about the non-equilibrium character of the process and the contribution of EP, since it has different analytical dependence on the cEQ and fEQ regimes, as discussed before in section˜2.2 and later in section˜5.
@@ 3.5 Comparison and differences to model-specific known results
Finite-time optimal protocols with kinks were analytically computed for a particle in a harmonic trap [10] and recently unified using optimal transport theory [11]. These model-specific results differ from the generalized finite-time framework in three aspects: (i) they optimize the free energy $\psi_{E}$ using control parameters of the equilibrium energy functional $E$, a close-to-equilibrium methodology (ii) the boundary conditions used to obtain kinks differ, and (iii) the thermodynamic cost of kinks is undefined in Ref.[10, 11], leading to the physical interpretation of kinks as a boundary artefact.
@@
We briefly summarize the procedure outlined in [11]. The boundary conditions used in [10, 11] are equivalent to fixing $A_{\alpha}^{i*}=A_{i}$ and computing $A_{\alpha}^{*int}\in(A_{\alpha}^{i},A_{\alpha}^{f})$ that minimize the bulk driving EP under the final value constraint. Thus, $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ is termed as the feasible driving protocol for the given finite-time $\tau$. Then, $A_{\alpha}^{*int}$ is used to compute $A_{\alpha}^{*i}$ and $A_{\alpha}^{*f}$ by introducing counter-diabatic driving terms that shifts the initially computed driving protocol $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ to $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ in driving time $\tau$, and it generates kinks of amplitude $A_{\alpha}^{i}-A_{\alpha}^{*i}$ and $A_{\alpha}^{f}-A_{\alpha}^{f*}$ in the optimal protocol. This implies that the computation of the initial geodesic for the finite-time process is based on the assumption of the absence of kinks at the initial time $t=0$. However, the finite-time optimal protocol $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ is obtained by shifting this protocol (counter-diabatic driving) to obtain the kink at both $t=0^{+}$ and $t=\tau^{-}$.
@@
This contradicts its formulation of constrained and unconstrained optimization for the initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, respectively. Hence, mathematically, this is interpreted as a boundary condition artifact because of the inability of assigning a thermodynamic cost to kinks within their framework. In contrast, within MinAP formalism, a thermodynamic cost is inherently assigned for kinks, and hence unconstrained optimization is implemented for both initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, signified by a double optimization problem in eq.˜17. Therefore, our optimization framework uses minimum assumptions, equivalently assuming the least information about optimal protocols. This allows us to implement an unconstrained optimization problem in eq.˜17 that searches/accesses a wider control parameter space for the optimal protocol.
@@
These differences lead to rather drastic physical results. For example, the short-time optimal protocol eq.˜23 is midpoint interpolation in the geodesic space $\mathcal{G}(A_{\alpha})$ and not necessarily in control parameter space $A_{\alpha}$ unless $\mathcal{G}(A_{\alpha})=A_{\alpha}$. Hence, even for short driving time $\tau$, the finite-time optimal protocol eq.˜23 respects the Riemannian geometric structure in control parameter space. In contrast, the short-time protocols obtained in Ref. [10, 11] follow a midpoint interpolation in control parameter space. This difference is irrelevant in the case of the optimal control of the trap centre, since $\mathcal{G}(A_{\alpha})=A_{\alpha}$ for the control of the trap centre and the short-time optimal protocols obtained in Ref. [81, 82]. However, $\mathcal{G}(A_{\alpha})\neq A_{\alpha}$ for the optimal control of the trap stiffness [24], giving divergent analytical results compared to Ref. [10, 11, 81, 82]. Subsequently, we will revisit this issue in section˜5.2 with the example of a stochastic particle in a harmonic trap and reconcile the discrepancies between optimal transport theory [49, 50, 58] and thermodynamic geometry [24].